Stock XYZ has an expected rate of return of 8% and risk of B= 0.40. Stock ABC has an expected return of 15% and B=1.30. The markets return is 12% and the risk free rate is 4%

A) According to CAPM, Which stock is a better buy?

B) What is the alpha on each stock? Plot the SML and each and each stocks risk-return point on one graph. Label everything appropriately and show the alphas graphically
Answer #1

Required Rate of XYZ = 0.04 + 0.40(0.12 - 0.04)

Required Rate of XYZ = 7.20%

Required Rate of ABC = 0.04 + 1.30(0.12 - 0.04)

Required Rate of ABC =14.40%

Stock XYZ is more undervalued so it should be bought.

Alpha of Stock XYZ = 0.08 - 0072 = 0.80%

Alpha of Stock ABC = 0.15 - 0.144 = 0.60%